# Local analysis of iterative reconstruction from discrete generalized Radon transform data in the plane

### v1

@@ 7. Computing the leading term $\Delta F_{0}$
In this section we consider $x=x_{0}+\epsilon\check{x}$ and assume for simplicity $x_{0}=0_{2}$. The reconstruction point is then $x=\epsilon\check{x}$. We partially return the dependence on $x$ in the notation by keeping only the dependence on $\check{x}$. If an argument of a function is $x_{0}=0_{2}$, then it is omitted from notation. Fix any $A_{0}>0$. Throughout this section we assume $|\check{x}|\leq A_{0}$. Define
@@
(7.3) $\begin{split}\Delta G(r,\check{t},\check{q};\theta,\epsilon):=&\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon\check{w})\frac{\Delta H(r,\check{w}+\check{t},\check{q};\theta,\epsilon)}{\check{w}}\text{d}\check{w},\\ \Delta H(r,\check{t},\check{q};\cdot):=&H(r+\check{t},\check{q};\cdot)-H(\check{t},\check{q};\cdot).\end{split}$
We define $\Delta G$ not via $G$ in (6.14) but in terms of $\Delta H$ because this allows for more accurate bounds. By (6.15), the expression for $\Delta F_{\epsilon}^{(l)}$ becomes:
(7.4) $\begin{split}&\Delta F_{\epsilon}^{(l)}(\epsilon\check{x})=-\frac{\epsilon^{1/2}}{\pi}\sum_{j_{1}}\Delta G\bigg{(}\frac{Q(\epsilon\check{x},\alpha_{j_{1}})-Q(\alpha_{j_{1}})}{\epsilon},\frac{Q(\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}.\end{split}$
As before, the sum is over $\alpha_{j_{1}}\in\text{dom}\,P$.
@@
Suppose assumptions 3.13.4 are satisfied. Pick any $A>0$. One has
(7.5) $\begin{split}\Delta G(r,\check{t},\check{q};\theta,\epsilon),\partial_{r}\Delta G(r,\check{t},\check{q};\theta,\epsilon)=O(|\check{t}|^{-1}),\ \check{t}\to\infty.\end{split}$
Also,
[…]
In the above formulas the big-$O$ terms are uniform in $\check{q}\in\mathbb{R}$, $\theta\in\text{dom}\,P$, $|\check{t}|\leq A/\epsilon$, $|r|\leq A$, and $0<\epsilon\ll 1$.
Recall that $Q(\alpha)$ stands for $Q(x_{0}=0_{2},\alpha)$. Suppose first that the selected $P(\alpha)$ is associated with $x_{0}$, so $Q(0)=0$. By Lemma D.3, $Q_{\alpha}^{\prime}(0)=0$. From
(7.7) $\begin{split}Q(\epsilon\check{x},\alpha)=&Q_{x}^{\prime}(0)\epsilon\check{x}+Q_{\alpha\alpha}^{\prime\prime}(0)\alpha^{2}/2+O(\epsilon^{3/2}),\\ Q_{\alpha}^{\prime}(\epsilon\check{x},\alpha)=&Q_{\alpha\alpha}^{\prime\prime}(0)\alpha+O(\epsilon),\quad\alpha=O(\epsilon^{1/2}),\end{split}$
we conclude
(7.8) $\alpha_{1}(\epsilon\check{x})=O(\epsilon^{1/2}),\ \alpha_{2}(\epsilon\check{x})=O(\epsilon),$
where $\alpha_{1}$, $\alpha_{2}$ are found by solving $Q(\epsilon\check{x},\alpha_{1})=0$ and $Q_{\alpha}^{\prime}(\epsilon\check{x},\alpha_{2})=0$. By (6.7), $Q_{x}^{\prime}(x,\alpha)=\Phi_{x}^{\prime}(x,\alpha)$.
@@
(7.9) $Q(\theta)\geq c\theta^{2},\quad|\theta|\leq\delta,$
for some $0<\delta\ll 1$. Also, $Q(\theta)\geq c$, $|\theta|\geq\delta$, $\theta\in\text{dom}\,P$. By (7.5), this implies
(7.10) $\begin{split}&\epsilon^{1/2}\bigg{(}\sum_{A\epsilon^{1/2}\leq|\alpha_{j_{1}}|\leq\delta}+\sum_{|\alpha_{j_{1}}|>\delta}\bigg{)}|\Delta G|\\ &\leq c\epsilon^{1/2}\bigg{(}\sum_{j_{1}>A\epsilon^{-1/2}}(\epsilon j_{1}^{2})^{-1}+\sum_{j_{1}=\delta/\epsilon}^{1/\epsilon}\epsilon\bigg{)}=O(A^{-1})+O(\epsilon^{1/2}),\end{split}$
where the arguments of $\Delta G$ are the same as in (7.4).
By (7.5) and (7.6), for $|\alpha_{j_{1}}|\leq A\epsilon^{1/2}$:
(7.11) $\begin{split}&\Delta G\bigg{(}\frac{Q(\epsilon\check{x},\alpha_{j_{1}})-Q(\alpha_{j_{1}})}{\epsilon},\frac{Q(\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}\\ &=\Delta G\bigg{(}\Phi_{x}^{\prime}\check{x}+O(\epsilon^{1/2}),\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2}+O(\epsilon^{1/2}),\frac{P(\epsilon\mu j_{1})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}\\ &=\Delta G\bigg{(}\Phi_{x}^{\prime}\check{x},\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}+\frac{O(\epsilon^{1/2}\ln(1/\epsilon))}{(1+\epsilon j_{1}^{2})^{3/2}}.\end{split}$
Pick any $P(\alpha)$, which is not associated with a segment of $\mathcal{S}$ through $x_{0}$. This means there is a neighborhood $U^{\prime}$ of $x_{0}$ such that none of the curves $\mathcal{S}_{(\alpha,P(\alpha))}$, $\alpha\in\text{dom}\,P$, is tangent to $\mathcal{S}\cap U^{\prime}$. There are two possibilities: (a) Neither of these curves passes through $x_{0}$ (in this case $Q(\alpha)$ is bounded away from zero on $\text{dom}\,P$) and (b) one of the curves contains $x_{0}$ (in this case $Q(\alpha)=0$ for some $\alpha\in\text{dom}\,P$).
If $|Q(\alpha)|>c$, $\alpha\in\text{dom}\,P$, then, similarly to (7.10),
(7.12) $\epsilon^{1/2}\textstyle\sum_{j_{1}}|\Delta G|\leq c\epsilon^{1/2}\textstyle\sum_{j_{1}=1}^{1/\epsilon}\epsilon=O(\epsilon^{1/2}).$
If $Q(\tilde{\alpha})=0$, but $|Q^{\prime}(\alpha)|>c$, then $|Q(\tilde{\alpha})|\asymp|\alpha-\tilde{\alpha}|$, $\alpha\in\text{dom}\,P$, and
(7.13) $\epsilon^{1/2}\textstyle\sum_{j_{1}}|\Delta G|\leq c\epsilon^{1/2}\ln(1/\epsilon)\textstyle\sum_{j_{1}=1}^{1/\epsilon}(1/j_{1})=O(\epsilon^{1/2}\ln^{2}(1/\epsilon)).$
Combine (7.10)–(7.13) and use (7.4)
(7.14) $\begin{split}\Delta F_{\epsilon}^{(l)}(\epsilon\check{x})=&-\frac{\epsilon^{1/2}}{\pi}\sum_{|\alpha_{j_{1}}|\leq A\epsilon^{1/2}}\Delta G\bigg{(}\Phi_{x}^{\prime}\check{x},\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}\\ &+O(\epsilon^{1/2}\ln^{2}(1/\epsilon))+O(1/A),\end{split}$
where $O(\epsilon^{1/2}\ln^{2}(1/\epsilon))$ is independent of $A$, and $P$ is associated with $x_{0}$. Here we have used that $\epsilon^{1/2}\sum_{j_{1}}(1+\epsilon j_{1}^{2})^{-3/2}<c$, $10<\epsilon<$ 1.
@@
(7.15) $\begin{split}\Delta F_{0}(\check{x})=-\frac{1}{\pi}\int_{|s|\leq A}\int_{0}^{1}&\big{[}G_{0}\big{(}\Phi_{x}^{\prime}\check{x}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}\\ &-G_{0}\big{(}(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}\big{]}\text{d}\check{q}\text{d}s+O(1/A),\end{split}$
@@
(7.16) $\begin{split}&G_{0}(\check{t},\check{q}):=\int_{\mathbb{R}}\frac{H_{0}(\check{w}+\check{t},\check{q})}{\check{w}}\text{d}\check{w},\\ &H_{0}(\check{t},\check{q}):=a_{0}(y_{0})\sum_{j_{2}}h_{0}^{\prime}(\check{t}+\check{q}-j_{2})(j_{2}-\check{q})_{+}^{1/2},\\ &h_{0}(u):=K(x_{0},\alpha_{0})\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(0)+u)\text{d}s.\end{split}$
@@
(7.17) $\begin{split}\Delta F_{0}(\check{x})=&-\frac{1}{\pi}\int_{\mathbb{R}}\int_{0}^{1}\big{[}G_{0}\big{(}\Phi_{x}^{\prime}\check{x}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}-G_{0}\big{(}(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}\big{]}\text{d}\check{q}\text{d}s\\ =&-\frac{a_{0}(y_{0})}{\pi}\int_{\mathbb{R}}\int_{\mathbb{R}}\int_{0}^{\infty}\frac{1}{\check{w}}\big{[}h_{0}^{\prime}\big{(}\check{w}+\Phi_{x}^{\prime}\check{x}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2}-\check{q}\big{)}\\ &\hskip 99.58464pt-h_{0}^{\prime}\big{(}\check{w}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2}-\check{q}\big{)}\big{]}\check{q}^{1/2}\text{d}\check{q}\text{d}s\text{d}\check{w}.\end{split}$
It is easy to see that $G_{0}(\check{t}+r,\check{q})-G_{0}(\check{t},\check{q})$ satisfies the same estimates as $\Delta G$ in (7.5), so the integral in (7.17) is absolutely convergent.
@@
(7.18) $\Delta F_{0}(\check{x})=C\int_{0}^{\Phi_{x}^{\prime}\check{x}}h_{0}(-t)\text{d}t,\ C:=\frac{a_{0}(y_{0})}{4(Q_{\alpha\alpha}^{\prime\prime}/2)^{1/2}}.$
@@
Pick any $A_{0}>0$. Throughout this section we assume $|\check{x}|\leq A_{0}$.

### v2

@@ 7. Computing the leading term $\Delta F_{0}$
In this section we consider $x=x_{0}+\epsilon\check{x}$ and assume for simplicity $x_{0}=0_{2}$. The reconstruction point is then $x=\epsilon\check{x}$. Fix any $A_{1}>0$. Throughout this section we assume $|\check{x}|<A_{1}$, including in Lemmas 7.2, 7.3.
From the proof of Lemma 6.2 and (6.13), (6.14), (6.17) it follows that
(7.1) $G(\check{t},\check{q};\epsilon\check{x},\theta,\epsilon)=G(\check{t},\check{q};0_{2},\theta,\epsilon)+O(\epsilon).$
By (6.18),
(7.2) $\begin{split}F_{\epsilon}^{(l)}(\epsilon\check{x})=-\frac{\epsilon^{1/2}}{\pi}\sum_{j_{1}\in J_{\alpha}}G\bigg{(}\frac{Q(x,\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon};0_{2},\alpha_{j_{1}},\epsilon\bigg{)}+O(\epsilon^{1/2}).\end{split}$
Define
@@
From (7.2) it follows that when computing $\Delta F_{0}(\check{x})$, we can replace $\epsilon\check{x}$ with $0_{2}$ in the list of arguments of $G$ when representing $F_{\epsilon}^{(l)}(\epsilon\check{x})$ in terms of $G$ in (7.4). Therefore, we can drop $x$ from the list of arguments of $h$, $H$, and $G$ and change the meaning of $*=(x,\theta,\epsilon)$ to $*=(\theta,\epsilon)$. Also, if the first argument of $Q$ is $x=0_{2}$, then it is omitted from notation.
@@
(7.5) $\begin{split}\Delta G(r,\check{t},\check{q};*):=&G(r+\check{t},\check{q};*)-G(\check{t},\check{q};*)\\ =&\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon\check{w})\frac{\Delta H(r,\check{w}+\check{t},\check{q};*)}{\check{w}}\text{d}\check{w},\\ \Delta H(r,\check{t},\check{q};*):=&H(r+\check{t},\check{q};*)-H(\check{t},\check{q};*).\end{split}$
We underscore that all the functions introduced in (7.3)–(7.5) are computed using a local patch of data $g=\chi_{V}\hat{f}$ as described in Lemma 4.2.
By (6.18), the expression for $\Delta F_{\epsilon}^{(l)}$ becomes:
(7.6) $\begin{split}&\Delta F_{\epsilon}^{(l)}(\epsilon\check{x})\\ &=-\frac{\epsilon^{1/2}}{\pi}\sum_{j_{1}\in J_{\alpha}}\Delta G\bigg{(}\frac{Q(\epsilon\check{x},\alpha_{j_{1}})-Q(\alpha_{j_{1}})}{\epsilon},\frac{Q(\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}.\end{split}$
@@
Suppose Assumptions 3.13.4 are satisfied. Pick any $A_{2}>0$. One has:
(7.7) $\begin{split}|\Delta G(r,\check{t},\check{q};*)|\leq c\frac{|r|\ln(1/|r|)}{1+|\check{t}|},\end{split}$
(7.8) $\begin{split}|\Delta G(r+h,\check{t},\check{q};*)-\Delta G(r,\check{t},\check{q};*)|\leq c\frac{|h|\ln(1/|h|)}{1+|\check{t}|},\end{split}$
and
(7.9) $|\Delta G(r,\check{t}+h,\check{q};*)-\Delta G(r,\check{t},\check{q};*)|\leq c\frac{|h|\ln(1/|h|)}{1+|\check{t}|}.$
In the above formulas, the $c$’s are independent of $|r|\leq A_{2}$, $|\check{t}|\leq A_{2}/\epsilon$, $|\check{q}|\leq A_{2}/\epsilon$, $\theta\in I_{\alpha}$, $0<\epsilon\ll 1$, and $0<h\ll 1$.
We will apply Lemma 7.1 to (7.6), so we select $A_{2}$ as follows:
(7.10) $A_{2}=\max_{|\check{x}|\leq A_{1},\alpha_{j_{1}}\in I_{\alpha},0<\epsilon\ll 1}\max\bigg{(}\frac{|Q(\epsilon\check{x},\alpha_{j_{1}})-Q(\alpha_{j_{1}})|}{\epsilon},|Q(\alpha_{j_{1}})|,|P(\alpha_{j_{1}})|\bigg{)}.$
Recall that $Q(\alpha)$ stands for $Q(x_{0}=0_{2},\alpha)$. Since $\check{x}$ is confined to a bounded set and $Q$ is smooth, the right-hand side of (7.10) is bounded.
Suppose first that $P(\alpha)$ is not associated with $x_{0}$. This means there is a neighborhood $U^{\prime}$ of $x_{0}$ such that no curve $\mathcal{S}_{(\alpha,P(\alpha))}$, $\alpha\in I_{\alpha}$, is tangent to $\mathcal{S}\cap U^{\prime}$. There are two possibilities: (a) Neither of these curves passes through $x_{0}$ (in this case $Q(\alpha)$ is bounded away from zero on $I_{\alpha}$) and (b) one of the curves contains $x_{0}$ (in this case $Q(\tilde{\alpha})=0$ for some $\tilde{\alpha}\in I_{\alpha}$).
(a) If $|Q(\alpha)|>c$, $\alpha\in I_{\alpha}$, the inequality (7.7) implies
(7.11) $\epsilon^{1/2}\textstyle\sum_{j_{1}\in J_{\alpha}}|\Delta G|\leq c\epsilon^{1/2}\textstyle\sum_{j_{1}=1}^{1/\epsilon}\epsilon=O(\epsilon^{1/2}).$
Here and below, the arguments of $\Delta G$ are the same as in (7.6).
(b) If $Q(\tilde{\alpha})=0$, but $|Q^{\prime}(\alpha)|>c$ on $I_{\alpha}$, then $|Q(\alpha)|\asymp|\alpha-\tilde{\alpha}|$, $\alpha\in I_{\alpha}$, and (7.7) leads to
(7.12) ${\color[rgb]{0,0,0}\epsilon^{1/2}\textstyle\sum_{j_{1}\in J_{\alpha}}|\Delta G|\leq c\epsilon^{1/2}\textstyle\sum_{j_{1}=1}^{1/\epsilon}\frac{\epsilon}{\epsilon j_{1}}=O(\epsilon^{1/2}\ln(1/\epsilon)).}$
Suppose now $P(\alpha)$ is associated with $x_{0}$. Therefore, $I_{\alpha}$ is a small neighborhood of $\alpha_{0}$. Recall that we assumed $\alpha_{0}=0$. By the definition of the function $Q$ (see (6.10)), $Q(0)=0$. Applying Lemma D.3 with $\tilde{\alpha}=\alpha_{0}=0$ and $\tilde{x}=x_{0}=0_{2}$ gives $Q_{\alpha}^{\prime}(0)=0$.
@@
(7.13) $Q(\alpha)\geq c\alpha^{2},\quad|\alpha|\leq\delta,$
for some $0<\delta\ll 1$. Also, $Q(\alpha)\geq c$, $|\alpha|\geq\delta$, $\alpha\in I_{\alpha}$. By (7.7), this implies
(7.14) $\begin{split}&\epsilon^{1/2}\bigg{(}\sum_{A\epsilon^{1/2}\leq|\alpha_{j_{1}}|\leq\delta}+\sum_{j_{1}\in J_{\alpha},|\alpha_{j_{1}}|>\delta}\bigg{)}|\Delta G|\\ &\leq c\epsilon^{1/2}\bigg{(}\sum_{j_{1}>A\epsilon^{-1/2}}\frac{\epsilon}{(\epsilon j_{1})^{2}}+\sum_{j_{1}=\delta/\epsilon}^{1/\epsilon}\epsilon\bigg{)}=O(A^{-1})+O(\epsilon^{1/2}),\end{split}$
By (6.10), (7.8) and (7.9), for $|\alpha_{j_{1}}|\leq A\epsilon^{1/2}$:
(7.15) $\begin{split}&\Delta G\bigg{(}\frac{Q(\epsilon\check{x},\alpha_{j_{1}})-Q(\alpha_{j_{1}})}{\epsilon},\frac{Q(\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}\\ &=\Delta G\bigg{(}\text{d}_{x}\Phi\cdot\check{x}+O(\epsilon^{1/2}),\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2}+O(\epsilon^{1/2}),\frac{P(\epsilon\mu j_{1})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}\\ &=\Delta G\bigg{(}\text{d}_{x}\Phi\cdot\check{x},\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}+O(\epsilon^{1/2}\ln(1/\epsilon)).\end{split}$
Combine (7.11), (7.12), (7.14), (7.15), and use (7.6) to obtain
(7.16) $\begin{split}\Delta F_{\epsilon}^{(l)}(\epsilon\check{x})=&-\frac{\epsilon^{1/2}}{\pi}\sum_{|\alpha_{j_{1}}|\leq A\epsilon^{1/2}}\Delta G\bigg{(}\text{d}_{x}\Phi\cdot\check{x},\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}\\ &+O(\epsilon^{1/2}\ln(1/\epsilon))+O(1/A).\end{split}$
Here we have used that there are $O(\epsilon^{1/2})$ values of $j_{1}$ such that $|\alpha_{j_{1}}|\leq A\epsilon^{1/2}$.
Our analysis shows (see (7.11), (7.12)) that if $\Gamma_{l}$ is not associated with $x_{0}$, then the data in $V_{l}=I_{\alpha}^{l}\times I_{p}^{l}$ does not contribute to the DTB at $x_{0}$.
@@
(7.17) $\begin{split}\Delta F_{0}(\check{x})=-\frac{1}{\pi}\int_{|s|\leq A}\int_{0}^{1}&\big{[}G_{0}\big{(}\text{d}_{x}\Phi\cdot\check{x}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}\\ &-G_{0}\big{(}(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}\big{]}\text{d}\check{q}\text{d}s+O(1/A),\end{split}$
@@
(7.18) $\begin{split}&G_{0}(\check{t},\check{q}):=\int_{\mathbb{R}}\frac{H_{0}(\check{w}+\check{t},\check{q})}{\check{w}}\text{d}\check{w},\\ &H_{0}(\check{t},\check{q}):=a_{0}(y_{0})\sum_{j_{2}\in\mathbb{Z}}h_{0}^{\prime}(\check{t}+\check{q}-j_{2})(j_{2}-\check{q})_{+}^{1/2},\\ &h_{0}(u):=K(x_{0},\alpha_{0})\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(0)+u)\text{d}s.\end{split}$
@@
(7.19) $\begin{split}&\Delta F_{0}(\check{x})\\ &=-\frac{1}{\pi}\int_{\mathbb{R}}\int_{0}^{1}\big{[}G_{0}\big{(}\text{d}_{x}\Phi\cdot\check{x}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}-G_{0}\big{(}(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}\big{]}\text{d}\check{q}\text{d}s\\ &=-\frac{a_{0}(y_{0})}{\pi}\int_{\mathbb{R}}\int_{\mathbb{R}}\int_{0}^{\infty}\frac{1}{\check{w}}\big{[}h_{0}^{\prime}\big{(}\check{w}+\text{d}_{x}\Phi\cdot\check{x}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2}-\check{q}\big{)}\\ &\hskip 99.58464pt-h_{0}^{\prime}\big{(}\check{w}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2}-\check{q}\big{)}\big{]}\check{q}^{1/2}\text{d}\check{q}\text{d}s\text{d}\check{w}.\end{split}$
Comparing (6.14) with (7.18), we see that the functions $H$ and $H_{0}$ are analogous and they have the same asymptotic behavior as $\check{t}\to\infty$. More precisely, $H_{0}$ satisfies Lemma D.1 with $\epsilon=0$. To apply this lemma to obtain the asymptotics of $H_{0}$, we set $\epsilon=0$ and drop all the restrictions of the kind $\check{t}<A/\epsilon$. The same modifications should be made in the proof of Lemma D.1 in Appendix D.2.
Likewise, the functions $G$ of (6.17) and $G_{0}$ of (7.18) are analogous and they also have the same asymptotic behavior as $\check{t}\to\infty$. Repeating the argument in the proof of (7.7) (see Appendix E.1 from the beginning through the paragraph following (E.5)) and replacing $H$ with $H_{0}$, we see that $G_{0}(\check{t}+r,\check{q})-G_{0}(\check{t},\check{q})$ satisfies the same estimates as $\Delta G$ in (7.7), so the first integral in (7.19) is absolutely convergent.
@@
(7.20) $\Delta F_{0}(\check{x})=C\int_{0}^{\text{d}_{x}\Phi\cdot\check{x}}h_{0}(-t)\text{d}t,\ C:=\frac{a_{0}(y_{0})}{4(Q_{\alpha\alpha}^{\prime\prime}/2)^{1/2}}.$
@@
Pick any $A_{0}>0$. Consistent with Theorem 3.6, throughout this section we assume $|\check{x}|<A_{0}$. As established in the preceding section, it suffices to consider only the segment $\Gamma_{l}$ (and the corresponding $a_{0}(y),P(\alpha)$) that is associated with $x_{0}$.